OpenStudy (anonymous):

HELP! A golf back is struck so that it leaves the head of the club with a speed of 50.0ms at an angle of 30.0° to the ground. What is the maximum height reached by the golf ball? ASSUME AIR RESISTANCE IS NEGLIGIBLE.

OpenStudy (anonymous):

In order to find out the maximum height, we only need concern ourselves with vertical motion. We can use some simple geometry to find out the vertical component of velocity: So the vertical velocity is 25 m/s. Gravity will be the force that decelerates the ball on its upward flight. We can use a kinematic equation to solve for when the ball comes to a stop (reaches the top of its flight). Vf^2 = Vi^2 + 2AX We will take Vf to be 0 m/s, as this is at the top of the flight. A = g = -9.8 m/s^2. 0 = (25)^2 - (2)(9.8)X -625 = -19.6 X X = 31.89 m The max height of the ball will be 31.89 m.